OpenStudy (anonymous):

a problems #4 my entrance exam pb:

OpenStudy (anonymous):

equation?

OpenStudy (anonymous):



OpenStudy (anonymous):

when\[ f(x+y) = f(x)+f(y)+4xy\] and \[f(1)=4\] find \[f(20)\] P.S. as ever, if you cannot do it within 10 minutes, give me your medal;))

OpenStudy (anonymous):

be calm man

OpenStudy (anonymous):

fx+fy = fx +fy +4xy i really dont know this lol

OpenStudy (anonymous):

i go learn it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thought this would take you guy some time :P

OpenStudy (anonymous):

you know i remember the hard time i'll have to write down from f(1)to f(20) :(( hope you guy have a better solution

OpenStudy (anonymous):

Apparently f(2) = 12, f(4) = 40, f(8) = 144, f(16) = 83,232, and so finally f(20) = f(16 + 4) = 13,400,392

OpenStudy (anonymous):

huh? i dont understand. could you explain it more clearly? jemurray btw as far as i can recall i got just 3 digit number

OpenStudy (anonymous):

Oh, oops, I made a typo in my calculator :) Just a second...

OpenStudy (anonymous):

nononooooooooooooooooooooooooooooooooooooooooooo im hungry:(

OpenStudy (anonymous):

f(16) = 416, f(20) = f(4 + 16) = 712

OpenStudy (zarkon):

I get 840 \[f(n)=2n(n+1)\]

OpenStudy (phi):

Here's one way to do it f(x+y)=f(x)+f(y)+4xy f(20)= f(1+19)= f(1)+4*1*19 + f(19) = 4+4*19 + f(19) f(20)= 4*20 + f(19) continue the pattern f(20)= 4*(1+2+...+20)= 4*20*21/2= 840

OpenStudy (anonymous):

Here is how I did it, I computed \( f(1+1)=f(2)\) , then \( f(3)=f(2)+f(1) \) then \(f(5)\),\( f(10) \) and finally \( f(20)\).